# Impact of memory on clustering in spontaneous particle aggregation

### v2

@@ 5 Equilibrium patterns in the large population limit
where $\chi_{[0,R]}$ denotes the characteristic function of the interval $[0,R]$, i.e., the kernel $W=W(x)$ corresponds to the sampling radius $R>0$. Using the semi-implicit finite difference discretization for the space variable and first-order forward Euler method for the time variable, the steady states are plotted for $R\in\{0.1,0.05,0.025\}$ in Figure 2. Let us note that the steady states are not unique and the particular pattern produced by the simulation depends on the choice of the initial datum; for all three simulations presented in Figure 2 we used $\varrho(t=0,x)=1+10^{-1}\sin(2\pi x(2-x))^{2}$. The same approach can be applied in higher-dimensional settings as well, see [BHW:2012:PhysD, Section 5.4] for the results with $d=2$.
Finally, let us mention that in the formal limit as $R\to 0+$ in (5.6) we obtain $W(x)=\delta(x)$, the Dirac delta function in $\mathbb{R}^{d}$. Then $W\ast\varrho=\varrho$ and equation (2.4) becomes a nonlinear but local reaction-diffusion equation. However, it does not produce any nontrivial patterns, since the following alternative holds: either $G=G(s)$ is such that all homogeneous steady states are asymptotically stable, or equation (2.4) is ill-posed. We refer to [BHW:2012:PhysD, Section 4.4] for more details.
@@ 6 Spontaneous aggregation in one spatial dimension ($d=1$)
For identification of the clusters we use the MATLAB implementation of the Density-based spatial clustering (DBSCAN) method [DBSCAN]. As we expect the clusters to be of size comparable to the sampling radius $R=0.025$, we set the parameter $\mathtt{epsilon}$, specifying the radius of a neighbourhood with respect to some point of the DBSCAN method, to $\mathtt{epsilon}=0.025$. Moreover, based on experimentation, we found that the choice of the second parameter of DBSCAN, $\mathtt{minPts}=20$, leads to the best results in identification of clusters. Examples of the results, recorded at the final timestep of the simulations with $K\in\{1,2,3,4,5\}$, are plotted in Figure 3. For comparison, we also simulate the system (2.1)–(2.3), i.e., spontaneous aggregation without memory; we refer to the corresponding results by $K=0$.
@@ 7 Spontaneous aggregation in two spatial dimensions ($d=2$)
where we choose the sampling radius $R=1/20=0.05$. We again discretize (3.1) using the Euler-Maruyama scheme with timestep $\Delta t=10^{-3}$. We initialize the simulation by generating initial positions ${\mathbf{x}}_{i}\in\Omega$, for $i\in[N]$, using a uniform distribution in $\Omega$, while all internal variables are initially equal to zero, i.e., ${\mathbf{y}}_{i}^{k}(0)={\mathbf{0}}$ for all $i\in[N]$ and $k\in[K]$. We again calculate $100$ stochastic realizations, now for $K\in\{1,2,\dots,8\}$, each time performing $10^{7}$ timesteps. We record the particle locations at the last timestep ($t=10^{4}$) and use these for evaluating clustering properties of the model. For identification of the clusters we again use the MATLAB implementation of the DBSCAN method [DBSCAN] with its parameter chosen as $\mathtt{epsilon}=0.05$ (same as the sampling radius) and $\mathtt{minPts}=12$.
@@ 8 Discussion
In this paper, we have shown that short-term memory enhances and long-term memory inhibits spontaneous particle aggregation. Our investigation has been based on the (first-order) spontaneous aggregation model without memory which has been previously investigated in the literature [BHW:2012:PhysD]. Its main properties are summarized in Section 2. The memory has been added into this model in Section 3 by introducing a chain of $K$ internal variables that allow the agents to ‘remember’ the densities they encountered in the past. If $K=1$, then our individual-based model is equivalent to what is called ‘the second-order model’ in reference [BHW:2012:PhysD], where the internal variable represents the agent’s velocity. Considering $K>1$, the model introduces additional layers of memory described by $K$ internal variables. Additional internal variables can be used to better fit the properties of relatively complex (high-dimensional) models of interacting particle systems, while keeping the number of degrees of the system (internal variables) relatively small [Erban:2016:CAM, Erban:2020:CGM]. For simplicity, our transfer of information between layers of memory is linear, but nonlinear functions can also be introduced to better fit the properties of some systems [Erban:2020:CGM, Gurney:2018:INN, Erban:2025:arxiv].
@@
We have performed systematic numerical simulations to evaluate the impact of the number of internal variables (or ‘memory layers’) $K$ on the dynamics of the many particle system, because (exact) analytical mathematical results to study these properties are not available and the development of the corresponding mathematical theories includes a number of open questions. Some progress in this direction can be made as shown in Section 4, where we have derived the formal macroscopic limit of the system as the number of agents tends to infinity, under the usual molecular chaos assumption [BHW:2012:PhysD, Erban:2012:ICB]. The limit is described by a Fokker-Planck equation, and we have characterized its steady states to gain an insight into the patterns (clusters) formed by the system.
Many organisms benefit from learning to adapt their behaviour to the distribution of resources and their learning depends on the memory capacity of each individual [Falcon:2023:LMS]. In our investigation, the number of memory layers, $K$, has been fixed, but it could be made variable for each agent. If the evolutionary objective is to develop large clusters in two-dimensional setting, then the choice $K=4$ can be considered ‘optimal’. Our memory model with $K$ layers could also be further extended by considering nonlinear activation functions to model a more complicated neural network [Gurney:2018:INN] and the applicability of the model could be further enhanced by introducing different modes of learning during the individual-based simulations [Lewis:2021:LAM].
In our computational simulations, we have used exponentially decaying function $G(s)=e^{-s}$ in equations (5.6) and (7.1) in order for clusters to be formed on effectively ‘intermediate’ timescales of the order $10^{3}$. Since all of the simulated particles are subject to the Brownian motion (including the particles in the clusters), the clusters are also able to slowly move (diffuse) on much longer timescales. However, as we are interested in biological applications, such long time behaviour of the model can be irrelevant, because other processes (e.g birth or death of individuals) could also happen on long timescales. We have therefore run the simulation until clusters emerge (starting from the initial random distribution of particles) and become ‘quasi-stationary’ in the sense that they do not move (on this timescale) and are in dynamic equilibrium with the outliers. For a study of the clustering behaviour on long timescales, see, e.g., [Carrillo:2019:AP].

### v3

@@ 5 Equilibrium patterns in the large population limit
where $\chi_{[0,R]}$ denotes the characteristic function of the interval $[0,R]$, i.e., the kernel $W=W(x)$ corresponds to the sampling radius $R>0$. Using the semi-implicit finite difference discretization for the space variable and first-order forward Euler method for the time variable, the steady states are plotted for $R\in\{0.1,0.05,0.025\}$ in Figure 2. Let us note that the steady states are not unique and the particular pattern produced by the simulation depends on the choice of the initial datum; for all three simulations presented in Figure 2 we used $\varrho(t=0,x)=1+10^{-1}\sin(2\pi x(2-x))^{2}$. The same approach can be applied in higher-dimensional settings as well, see [8, Section 5.4] for the results with $d=2$.
Finally, let us mention that in the formal limit as $R\to 0+$ in (5.6) we obtain $W(x)=\delta(x)$, the Dirac delta function in $\mathbb{R}^{d}$. Then $W\ast\varrho=\varrho$ and equation (2.4) becomes a nonlinear but local reaction-diffusion equation. However, it does not produce any nontrivial patterns, since the following alternative holds: either $G=G(s)$ is such that all homogeneous steady states are asymptotically stable, or equation (2.4) is ill-posed. We refer to [8, Section 4.4] for more details.
@@ 6 Spontaneous aggregation in one spatial dimension ($d=1$)
For identification of the clusters we use the MATLAB implementation of the Density-based spatial clustering (DBSCAN) method [30]. As we expect the clusters to be of size comparable to the sampling radius $R=0.025$, we set the parameter $\mathtt{epsilon}$, specifying the radius of a neighbourhood with respect to some point of the DBSCAN method, to $\mathtt{epsilon}=0.025$. Moreover, based on experimentation, we found that the choice of the second parameter of DBSCAN, $\mathtt{minPts}=20$, leads to the best results in identification of clusters. Examples of the results, recorded at the final timestep of the simulations with $K\in\{1,2,3,4,5\}$, are plotted in Figure 3. For comparison, we also simulate the system (2.1)–(2.3), i.e., spontaneous aggregation without memory; we refer to the corresponding results by $K=0$.
@@ 7 Spontaneous aggregation in two spatial dimensions ($d=2$)
where we choose the sampling radius $R=1/20=0.05$. We again discretize (3.1) using the Euler-Maruyama scheme with timestep $\Delta t=10^{-3}$. We initialize the simulation by generating initial positions ${\mathbf{x}}_{i}\in\Omega$, for $i\in[N]$, using a uniform distribution in $\Omega$, while all internal variables are initially equal to zero, i.e., ${\mathbf{y}}_{i}^{k}(0)={\mathbf{0}}$ for all $i\in[N]$ and $k\in[K]$. We again calculate $100$ stochastic realizations, now for $K\in\{1,2,\dots,8\}$, each time performing $10^{7}$ timesteps. We record the particle locations at the last timestep ($t=10^{4}$) and use these for evaluating clustering properties of the model. For identification of the clusters we again use the MATLAB implementation of the DBSCAN method [30] with its parameter chosen as $\mathtt{epsilon}=0.05$ (same as the sampling radius) and $\mathtt{minPts}=12$.
@@ 8 Discussion
In this paper, we have shown that short-term memory enhances and long-term memory inhibits spontaneous particle aggregation. Our investigation has been based on the (first-order) spontaneous aggregation model without memory which has been previously investigated in the literature [8]. Its main properties are summarized in Section 2. The memory has been added into this model in Section 3 by introducing a chain of $K$ internal variables that allow the agents to ‘remember’ the densities they encountered in the past. If $K=1$, then our individual-based model is equivalent to what is called ‘the second-order model’ in reference [8], where the internal variable represents the agent’s velocity. Considering $K>1$, the model introduces additional layers of memory described by $K$ internal variables. Additional internal variables can be used to better fit the properties of relatively complex (high-dimensional) models of interacting particle systems, while keeping the number of degrees of the system (internal variables) relatively small [27, 28]. For simplicity, our transfer of information between layers of memory is linear, but nonlinear functions can also be introduced to better fit the properties of some systems [28, 33, 29].
@@
We have performed systematic numerical simulations to evaluate the impact of the number of internal variables (or ‘memory layers’) $K$ on the dynamics of the many particle system, because (exact) analytical mathematical results to study these properties are not available and the development of the corresponding mathematical theories includes a number of open questions. Some progress in this direction can be made as shown in Section 4, where we have derived the formal macroscopic limit of the system as the number of agents tends to infinity, under the usual molecular chaos assumption [8, 23]. The limit is described by a Fokker-Planck equation, and we have characterized its steady states to gain an insight into the patterns (clusters) formed by the system.
Many organisms benefit from learning to adapt their behaviour to the distribution of resources and their learning depends on the memory capacity of each individual [32]. In our investigation, the number of memory layers, $K$, has been fixed, but it could be made variable for each agent. If the evolutionary objective is to develop large clusters in two-dimensional setting, then the choice $K=4$ can be considered ‘optimal’. Our memory model with $K$ layers could also be further extended by considering nonlinear activation functions to model a more complicated neural network [33] and the applicability of the model could be further enhanced by introducing different modes of learning during the individual-based simulations [43].
In our computational simulations, we have used exponentially decaying function $G(s)=e^{-s}$ in equations (5.6) and (7.1) in order for clusters to be formed on effectively ‘intermediate’ timescales of the order $10^{3}$. Since all of the simulated particles are subject to the Brownian motion (including the particles in the clusters), the clusters are also able to slowly move (diffuse) on much longer timescales. However, as we are interested in biological applications, such long time behaviour of the model can be irrelevant, because other processes (e.g birth or death of individuals) could also happen on long timescales. We have therefore run the simulation until clusters emerge (starting from the initial random distribution of particles) and become ‘quasi-stationary’ in the sense that they do not move (on this timescale) and are in dynamic equilibrium with the outliers. For a study of the clustering behaviour on long timescales, see, e.g., [9].